OpenStudy (anonymous):

Keith applies a new brand of fertilizer to his grass. He finds that the rate at which the grass grows can be modeled by a linear function. He observes that the rate of growth is 1 centimeter every 3 days. After 18 days, he measures the height of the grass to be 8 centimeters. The initial height of the grass was ______centimeters tall.

OpenStudy (anonymous):

medal and fan

Directrix (directrix):

Let h be the initial height of the grass Every 3 days, the grass grows 1 inch. The rate of change is 1/3 which represents a 1 inch increase in height per 3 days of time. y represents the height of the grass at time x days. y = h + (1/3) * x

Directrix (directrix):

Look at the point (18, 8) which means 18 days results in 8 inches of growth. y = h + (1/3) * x 8 = h + (1/3 * 18) Solve for h and you will get the initial grass height. @muneeb180 Post what you get, okay?

OpenStudy (anonymous):

lol gotta love the study island stuff.... im doing it late rn soo yea rushing to get them